OpenStudy (anonymous):

Write and solve a system of equations that represents each situation. Interpret the situation. Mr. Blackwell's company produces nuts and bolts. The total monthly profit for his company was $76,378. The profit earned from nuts was $3,428 more than the profit earned from bolts.

OpenStudy (ybarrap):

\( P=76,378\\ N=3,428+B\\ P=N+B\\ =3,428+B + B\\ =3,428+2B\\ =76,378\\ \implies 76,378-3,428=2B\\ \implies B=\cfrac{76,378-3,428}{2} \) Where P is total Profit, N is the profit from nuts and B is the profit from Bolts. Now you know the profits for Bolts and Nuts.